Write a two column proof:
If one angle in a linear pair is a right angle, then the other is also a right angle.
Given:
Prove:
7 statements and 7 reasons.
I need this now, please. Thanks! ASAP, I beg you.

Answer by jim_thompson5910(35256)

Basic idea:

A linear pair of angles are supplementary. So if x and y are a linear pair of angles, then x+y = 180

If one angle is a right angle, then x = 90. So

x+y = 180

90+y = 180

y = 180-90

y = 90

which means that the other angle is a right angle